kekeman:

(2x^2-4xy+2y^2)(3x-4) A.) 6x^3-12x^2y+6xy^2-8x^2+16xy-8y^2 B.) 6x^3-6x^2y-8x^2+16xy-8y^2 C.) 2x^2-4xy+2y^2+3x-4 D.) 6x^3-12x^2y+6xy^2-8x^2-16xy+8y^2

AZ:

Just distribute the terms! 2x^2 multiplied with 3x and -4 -4xy multiplied with 3x and -4 2y^2 multiplied with 3x and -4 (2x^2-4xy+2y^2)(3x-4) = (2x^2 * 3x) + (2x^2 * -4) + (-4xy * 3x) + (-4xy * -4) + (2y^2 * 3x) + (2y^2 * -4) and simplify

kekeman:

Expand by distributing sum groups. 2x^2(3x-4)-4xy(3x-4)+2y^2(3x-4) Expand by distributing terms. 6x^3-8x^2-4xy(3x-4)+2y^2(3x-4) Expand by distributing terms. 6x^3-8x^2-(12x^2y-16xy)+2y^2(3x-4) Expand by distributing terms. 6x^3-8x^2-(12x^2y-16xy)+6y^2x-8y^2 Remove parentheses. = 6x^3-8x^2-12x^2y+16xy+6y^2x-8y^2

kekeman:

@az wrote:
Just distribute the terms! 2x^2 multiplied with 3x and -4 -4xy multiplied with 3x and -4 2y^2 multiplied with 3x and -4 (2x^2-4xy+2y^2)(3x-4) = (2x^2 * 3x) + (2x^2 * -4) + (-4xy * 3x) + (-4xy * -4) + (2y^2 * 3x) + (2y^2 * -4) and simplify
Is that correct?

AZ:

That is correct! If you re-arrange the terms a little bit, you'll get 6x^3-12x^2y+6xy^2-8x^2+16xy-8y^2

kekeman:

@az wrote:
That is correct! If you re-arrange the terms a little bit, you'll get 6x^3-12x^2y+6xy^2-8x^2+16xy-8y^2
Which is A.)

AZ:

Exactly!

kekeman:

Ok thanks for the help!!!

AZ:

It was my pleasure!